OpenStudy (anonymous):

Help me convert to rectangular coordinates (-7,(2 pi)/3)

OpenStudy (anonymous):

@jdoe0001 could you help me out?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \left(-7,\frac{2\pi}{3}\right)\\ \quad \uparrow \quad \uparrow \\ \quad r\quad \theta \end{array} recall\implies \begin{array}{llll} x=rsin(\theta)\\ y=rcos(\theta) \end{array}\)

OpenStudy (jdoe0001):

ehh rather \(\large \begin{array}{llll} \left(-7,\frac{2\pi}{3}\right)\\ \quad \uparrow \quad \uparrow \\ \quad r\quad \theta \end{array} recall\implies \begin{array}{llll} x=rcos(\theta)\\ y=rsin(\theta) \end{array} \\ \quad \\ \implies (x,y)\implies [rcos(\theta),rsin(\theta)]\)

OpenStudy (jdoe0001):

so check your Unit Circle for the values of that angle :)